Trace the curve (x/a)^2/3 + (x/a) ^2/3=1
1
Expert's answer
2019-09-11T09:34:57-0400

The curve equation can be transformed as follows:


"(\frac{x}{a})^{2/3}+(\frac{x}{a})^{2/3}=1""(\frac{x}{a})^{2/3}=\frac{1}{2}""\frac{x}{a}=(\frac{1}{2})^{3/2}""x=a(\frac{1}{2})^{3/2}"

This equation represents vertical line, at "x=\frac{a}{2^{3/2}}=const".

The line is vertical as it is independent of "y"


As an example, the plot for "a=1" is attached